# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ I Introduction
II Development of Cooperative Perception: From Early Time Exploration till Recent Research
@@
Figure 4: Overview of collaboration schemes for autonomous driving.
@@
Contemporary CP strategies encompass three collaboration styles: early, intermediate, and late collaboration, each characterized by the nature of shared data. To enhance the efficacy of multi-agent cooperation, certain studies have embraced hybrid collaboration approaches, as exemplified in [110], combining early and late collaboration to consolidate information from diverse infrastructures. Fig. 4 provides a schematic overview of these collaboration schemes, considering a scenario involving two CAVs.
As illustrated in Fig. 4, early collaboration involves aggregating raw sensor observations collected by all participating agents, fostering a comprehensive perspective [111, 112, 113]. This approach is capable of addressing challenges like occlusion and long-range obstacles encountered in single-agent perception. However, its practical implementation necessitates substantial communication bandwidth. In contrast, late collaboration involves exchanging perceptual outcomes, requiring minimal data transmission [114]. While efficient in communications resource usage, late collaboration may introduce inherent noise and incompleteness into individual perceptual outputs, potentially resulting in suboptimal fusion outcomes. In response to the trade-off between performance and bandwidth constraints, intermediate collaboration has emerged as a viable solution. This approach aggregates middle features across multiple agents, incorporating representative information within compressed features. This integration holds promise for enhancing both communication efficiency and perception capabilities.
@@
Traditional approaches in CP have typically relied on late cooperation strategies due to the complexity of feature extraction and limited communication capabilities. These strategies operate at the object level or track level. However, with the rapid progress of deep learning, intermediate collaboration has gained prominence in multi-agent cooperation, primarily due to its ability to alleviate the burden on communication resources [115]. Nevertheless, it risks introducing extraneous and unnecessary information, particularly in dynamic networking conditions. To address this challenge, optimizing cooperative information and collaborator selection becomes imperative within a CP system. Furthermore, deep representation introduces higher levels of abstraction and implicitness in the spatial interactions between agents, necessitating suitable data alignment and information fusion algorithms. Developing a unified CP framework to address these issues is of significant importance in advancing Cooperative Driving Automation (CDA).
III V2X Communication
V2X communication is the fundamental architecture that enables CP applications for autonomous vehicles. This section reviews recent advancements in V2X communications for CP applications, focusing on four enabling technologies: edge computing, blockchain, digital twins, and 6G communications.
III-A Edge Computing-enabled V2X Communications
In modern CP systems, addressing the computational and storage requirements of extensive perception data gathered from vehicular sensors and devices is essential. Cloud computing and data centres have emerged as prevalent solutions, utilizing V2X communications to interconnect vehicular devices and servers for data sharing. However, these solutions can introduce significant communication overhead and offer limited computational flexibility for real-time CP tasks. A more efficient alternative is edge computing, which positions data and computation services at the network’s edge, in close proximity to vehicular devices, facilitating rapid response in CP operations [116].
As an example, edge computing has been proposed for V2X communication links in [117] to address computation-intensive tasks and facilitate timely decisions for data offloading, a critical requirement in real-time CP scenarios due to its proximity to end users. A low-latency task offloading model was established, allowing for the offloading of CP multi-tasks via V2X uplinks, managed by a mobile edge computing server positioned at Roadside Units (RSUs). A mobility-aware dynamic offloading strategy was introduced to account for the dynamic nature of vehicles during task offloading, capturing time-varying characteristics such as vehicular mobile positions, V2X communication status, and data task specifics.
In a similar work, computation offloading was employed to enable rapid-response CP services by integrating V2X communications into edge computing [118]. Specifically, a novel joint optimization framework was presented, addressing the allocation of caching space and computation resources for vehicular edge servers. The aim was to minimize offloading latency by considering factors such as offloading decisions and edge computing capabilities, including maximum bandwidth resources and available data storage. Beyond data offloading, edge computing also provides a viable solution for low-latency content caching, with edge servers, including base stations and RSUs, employed to collect and process data offloaded by nearby vehicular devices in real-time [119]. In this context, efficient resource management, encompassing resource allocation for V2X communication channels and load balancing among edge servers, is pivotal in ensuring effective content caching for CP applications.
Clustering presents an intriguing approach to further mitigate communication latency in V2X links, where edge computing services are partitioned into distinct zones, with each zone exclusively serving a specific number of connected vehicles to enhance CP task efficiency [120]. In this setup, each CP task can be managed by an RSU master within its coverage area, while computing tasks from vehicles in other zones are routed to the appropriate RSUs to prevent communication interference. In a related study, efforts were made to minimize V2X communication latency in vehicular services, including CP, by jointly considering both 5G-enabled uplink and downlink delays, data retrieval overheads, and content execution latency [121]. Latency assessments were conducted under diverse vehicular 5G communication configurations, demonstrating the advantages of edge computing in terms of content execution time savings compared to alternatives like cloud computing.
Another investigation on latency for edge computing in V2X communication was introduced in [122], where a radio latency model was formulated to depict the interrelation between backhaul latency at the edge server and data transmission latency between the edge server and end users based on Signal-to-Noise Ratio measurements at the evolved NodeB. Simulation results unveiled the positive correlation of vehicle density, edge server location, and data execution delays, providing valuable insights into integrating edge computing services into V2X communications to minimize data content processing time.
Moving from a single-edge computing scenario, the work in [123] extended the concept to a multi-edge computing framework, where distributed edge servers collaboratively serve the shared network of end users in a three-layer V2X architecture. At the lowest layer, vehicles and end users can offload their data to execute CP algorithms hosted at the edge server, managed at the second layer. Given the intricate nature of edge computing resources, an energy-aware edge service selection algorithm was developed to choose the suitable edge server for service requests from end users. In extensive data processing cases, end users would offload to the cloud server at the third layer. To effectively handle various data requests, a task priority classification scheme was integrated to ensure equitable data execution among different end-user groups.
III-B Blockchain-enabled V2X Communications
In addition to latency considerations, security is paramount in information exchange and data sharing over V2X communications among CAVs and agents within CP systems. Inadequate data protection and security measures can leave communications vulnerable to attacks, potentially leading to information theft and posing critical challenges to the viability of CP applications [124]. Another vulnerability arises from the traditional CP networking topology, where a centralized server predominantly manages perception algorithms. A breach of this central server can result in the collapse of the entire CP system. Consequently, there is a pressing need for a more decentralized CP solution, with blockchain technology emerging as a promising candidate due to its secure and decentralized data ledger capabilities.
Blockchain offers unique security attributes, such as decentralization, availability, traceability, and trustworthiness, presenting intriguing solutions encompassing decentralized V2X communications, distributed collaborative information perception among CAVs, and information verification and user authentication for secure CP services [125]. For instance, a proof-of-concept was presented in [126], utilizing blockchain for decentralized V2X communications. Vehicular data and messages from CAVs are encrypted and stored as transactions within the blockchain, distinguishable by vehicle identification and public keys. Rather than relying on a central authority, blockchain facilitates the creation of a message management ecosystem where distributed vehicles possess equal rights to monitor and verify information through blockchain accounts embedded within their software. Smart contracts, self-executing software running atop the blockchain, were employed to establish a straightforward mechanism for decentralized governance and information authentication [127].
In [128], a Proof-of-Local Learning mechanism was introduced, encrypting and sharing data collected by CAVs through a chain of blocks monitored by a distributed network of RSUs. Blockchain facilitates equitable information exchange and low-latency CP information storage via edge computing. Blockchain technology was also harnessed in [129] to enhance secure V2X communications within a three-layer CP system, encompassing the perception, network, and blockchain layers. The perception layer includes an Android auto application for CAVs, offering parking space and vehicle localization services. The network layer facilitates interconnection between vehicles, RSUs, and the cloud server, while the blockchain layer delivers security services, including immutability, integrity, and confidentiality for information exchanged across the other two layers. Smart contracts were integrated to provide traceable payment for vehicles, offering a mobile transaction cost calculation mechanism that ensures equitable payment for data storage and retrieval requests. Furthermore, blockchain-based smart contracts were employed in [130] for V2X links, showcasing their capability to address jamming attacks aimed at intercepting messages shared over V2X links, even when encrypted. The tracing capability of blockchain technology enables tracking shared information, including its source and destination, through transaction logs. NS-3 simulator simulations showed that blockchain allows for low-latency information exchange and V2X communication in a decentralized networking environment.
Moreover, in a study by [131], a blockchain-based transmission solution for V2X communications was introduced, focusing on the block mining mechanism. Blockchain mining is essential for data verification but can be energy-intensive and result in significant communication overhead. To address this, an offloading methodology was proposed, enabling the offloading of mining tasks to mining clusters rich in computing resources dedicated to transaction verification. Meanwhile, perception messages among CAVs can be exchanged without incurring additional latency costs, aiming to deliver a high-quality user experience in CP applications. In addition to data protection, blockchain can offer intrusion detection capabilities to enhance the security of V2X communications [132]. A micro-blockchain scheme was devised, facilitating reliable intrusion detection through a joint design involving one macro-blockchain and several micro-blockchains. The quality of local intrusion samples, a key factor for efficient intrusion detection, is securely managed through a tamper-resistant ledger integrated into the vehicle network. However, when deploying blockchain for secure V2X communications in CP applications, addressing various critical technical challenges, including scalability, high transaction latency, and limited flexibility [133] stemming from complex ledger settings, is crucial. Comprehensive solutions need to be further explored to ensure robust blockchain-based V2X communications.
III-C Digital Twin-enabled V2X Communications
In V2X communications for CP, a large amount of data is collected from surrounding environments, including vehicles and RSUs. To enhance the efficiency of CP algorithms, digital reproduction of the physical V2X environment by involving such data would be an attractive solution where Digital Twin (DT) emerges as an ideal technology candidate. DT-empowered V2X networks allow the creation of a cyber 3D platform to represent a physical V2X system by combining multi-modal sensory data such as messages and signals from CAVs and CP agents.
For instance, a DT-based V2X system introduced in [134] integrated a 3D virtual urban environment and traffic simulation into an information perception system comprising distributed vehicles and a base station. To construct an urban mobility simulator, this system gathered perceptive traffic information, encompassing various elements such as stochastic vehicle types, traffic light systems, and automated driving behaviours. This DT-based solution reduces testing costs associated with on-road experiments. It provides a practical environment for evaluating the quality of V2X communications, including channel estimation, which can be implemented using neural networks. In a different context, a DT-based V2X communication system founded on vehicle-to-cloud computing was tailored for a CP-enabled cooperative ramp merging application [135]. Field implementation was conducted with three typical passenger vehicles, demonstrating that the proposed system enhances current ramp merging scenarios’ safety and environmental sustainability. A realistic DT model was constructed to manage computations at the cloud server, coordinating data sharing from distributed vehicles via 4G networks. A CP simulation system was established within this framework, enabling virtual drivers to control their vehicles with DT support, facilitating safe and smooth cooperative merging maneuvers within the same coverage area.
DT technology also found application in V2X communication visualization through multi-vehicle experiments [136], incorporating both physical and virtual vehicles to execute collaborative perception tasks. A DT model was introduced, featuring two designs: cyberspace with twin objects and physical space with a localization testbed. Cameras captured vehicle motion images, which were shared with the cloud server via V2X communications for motion analytics, including the position and heading of vehicles. Furthermore, DT plays a role in assessing the reliability of V2X communications. In [137], a site-specific DT framework was introduced to evaluate the error rate of vehicular wireless links through a combination of geometry-based channel model simulation and realistic hardware-in-the-loop experiments. Reflective objects such as traffic signs and vehicles were employed to simulate data transmission in both simulations and testbeds, demonstrating a strong correlation in transmission error rate estimation.
To cope with multi-modal V2X communications where vehicular data such as traffic signals and vehicle messages are multi-modal, a DT-based framework was introduced in [138] as high-definition 3D mapping for physical V2X communication environments. Ego vehicles are equipped with multiple sensors, such as cameras, inertial measurement units, and radars, to perceive information from surroundings. The DT model was developed through three stages: sensory data acquisition from sensors, semantic instance segmentation, and digital 3D reconstruction. Several evaluations were conducted on an unreal engine-based automotive simulator for the DT model, showing feasible V2X channel modelling and efficient sensory data visualization. This research is essential in providing further insights into using DT for high-quality V2X communication analytics and perceptive information processing. Another similar line of research is reference [139], which suggested a DT-based CAV communication model by considering V2X communications and driving behaviour patterns. This framework was useful for forming cooperation of distributed drivers, especially at non-signalized intersections, where a cooperative motion control solution was integrated to generate the vehicle speed trajectories to achieve traffic flow consensus concerning communication constraints, including delay and packet loss.
TABLE III: Summary of the Works of V2X Communications.
Category Method Year Scenario Contribution
Edge Computing [117] 2021 V2X An edge computing model was proposed for low-latency task offloading in V2X communications. A mobility-aware dynamic offloading strategy was introduced to handle the dynamics of CAVs during task offloading.
[118] 2019 V2V A computation offloading was proposed to enable fast-response CP services by integrating V2X communications into edge computing.
[121] 2019 V2X An optimization framework was developed to minimize V2X communication latency, including CP, in vehicular services by jointly considering 5G-enabled uplink and downlink delays, data retrieval overheads, and content execution latency.
[123] 2021 V2V A multiple-edge computing framework was presented in a V2X communication setting, where distributed edge servers are jointly employed to serve the shared network of end users in a three-layer V2X architecture.
Blockchain [126] 2021 V2X A proof-of-concept was proposed using blockchain for decentralized V2X communications. The vehicular data and messages from CAVs can be encrypted and stored in the blockchain as a ledger of transactions, which are distinguished by vehicle identification and public keys.
[128] 2022 V2V A proof-of-local learning mechanism was proposed where data collected by CAVs are encrypted and shared by a chain of blocks jointly monitored by a distributed network of RSUs.
[130] 2023 V2X A blockchain-based smart contract scheme was provided for V2X links. Blockchain was used to cope with jamming attacks, aiming to steal messages shared over V2X links via an immutable data ledger.
[132] 2019 V2V A micro-blockchain scheme was built to offer reliable intrusion detection enabled by a joint design of one macro-blockchain and several micro-blockchains.
Digital Twin [134] 2022 V2X A DT-based V2X system was proposed to integrate the 3D virtual urban environment and traffic simulation in an information perception system consisting of distributed vehicles and a base station.
[135] 2022 V2V A DT-based V2X communication system was considered using vehicle-to-cloud computing for a CP-enabled cooperative ramp merging application.
[136] 2022 V2X DT was investigated in V2X communication visualization through multi-vehicle experiments, using both physical and virtual vehicles to execute collaborative perception tasks.
[137] 2022 V2X A DT-based CAV communication model was proposed by taking into account V2X communications and driving behaviour patterns.
6G [140] 2022 V2X A trial using 6G for V2X communications was introduced where intelligent reflecting surfaces were integrated to facilitate V2X communications by addressing signal blockage and large-scale fading due to other vehicles and ubiquitous objects in urban cities.
[141] 2023 V2X 6G-based intelligent reflecting surfaces were exploited to optimize the achievable transmission rate for V2X communication links between the base station and CAVs.
[142] 2023 V2X A resource allocation strategy for 6G-based V2X communications was studied based on reinforcement learning, which enabled the execution of allocation decisions in an environment created by the interaction of the base station and end users.
[143] 2022 V2X A cooperative collision detection mechanism was proposed, where reinforcement learning was applied to establish a near-optimal policy for vehicle trajectories and V2X communications in a 6G wireless network setting.
III-D 6G-enabled V2X Communications
Enabled by the revolution of wireless communication technology with 6G as an emerging wireless generation, V2X communications in CP applications are envisioned to be significantly upgraded with many new interesting features offered by 6G [86]. These would be ultra-low communication latency, high data rate, and connection reliability, with improved energy and spectrum efficiency compared to its previous wireless generations, such as 5G [144]. A trial using 6G for V2X communications was introduced in [140], where intelligent reflecting surfaces were integrated to facilitate V2X communications by addressing signal blockage and large-scale fading due to other vehicles and ubiquitous objects in urban cities. Thus, intelligent reflecting surfaces were employed as signal relays to assist data transmissions, aiming to achieve two performance targets: latency reduction and transmission rate enhancement [145]. 6G-based intelligent reflecting surfaces were also exploited to optimize the achievable transmission rate for V2X communication links between the base station and CAVs [141]. Channel gains for vehicle-based station communications were formulated, phase-shift matrices were constructed in a fashion, and the data rate was optimized via a simplified convex optimization method. In addition to transmission rate optimization for V2X Communications, 6G offers flexible channel management solutions via Multiple-input Multiple-output (MIMO) settings, where beamforming management plays a crucial role [146]. In this regard, a channel subspace structure was constructed using received vehicle signals evaluated via perceptive ray-tracing channel data and vehicle trajectories.
Integrating 6G technology into V2X communications plays a pivotal role in ensuring reliable connectivity, a critical requirement for CP applications where timely sharing of sensory data from vehicles with servers, such as base stations, is essential. Effective resource management is paramount in guaranteeing adequate resources to all vehicles to facilitate successful and swift data transmission. This incentive led the work in [142] to propose a resource allocation strategy tailored for 6G-based V2X communications. This strategy allocates resource blocks to participating CAVs within a shared vehicular network, guided by reinforcement learning principles. This approach allows for formulating allocation decisions within an environment shaped by interactions between the base station and end users. Unlike conventional optimization methods that rely on complex mathematical formulations, reinforcement learning estimates resource allocation in an equitable and optimal manner while mitigating transmission interference, all within the constraints of radio resources.
Figure 5: Illustration of a 6G-based V2X communication system.
Furthermore, one of the significant challenges in CP applications revolves around potential conflicts among vehicles, particularly when they share the same wireless channels for data transmission and operate within the same coverage area. The work in [143] proposed a cooperative collision detection mechanism to address this issue. This mechanism leverages reinforcement learning techniques to establish a near-optimal policy governing vehicle trajectories and V2X communications in the context of a 6G wireless network environment. The optimization process considers factors such as latency and quality of service in accordance with 6G standards, utilizing practical datasets to evaluate collision possibilities in dense vehicle networks.
For convenience, the related works in V2X communications are summarized in Table III, highlighting each reference work’s key contribution. Fig. 5 shows an example of 6G-based V2X communication systems. Regarding the standardization of V2X communication, protocol settings and corresponding use cases, the readers can refer to [147].
@@ IV-A Framework Overview
The collaboration stage encompasses three primary components, as depicted in Fig. 6: preparation for transmission, cooperative information alignment, and perception information fusion. To ensure efficient communication while preserving the accuracy of the cooperation system, the initial step involves carefully selecting suitable cooperative agents and pertinent perceptual information. Subsequently, each agent shares the CAM and the refined CPM with its peers through V2X communication. Following this, the ego vehicle aligns the acquired perception information with local data through coordinate conversion or pose transformation techniques. Finally, the aggregated data undergoes consolidation via a fusion algorithm, followed by decoding, resulting in a conclusive perception outcome. Note that all CAVs can transmit their perceptual data to a roadside infrastructure via V2X networks. The associated edge computing infrastructure is responsible for aggregating and consolidating this incoming data, which is then distributed to CAVs according to their specific needs. In the subsequent sections, we provide detailed insights and a review of existing works on the collaboration stage’s primary components.
IV-B Preparation for Transmission
TABLE IV: Summary of the Methods of Preparation for Transmission.
@@ IV-B Preparation for Transmission
CAS CPOB-LearnCom [150] 2022 V2I LiDAR CARLA-3D [150] Intermediate Detection $\times$
CODS [151] 2023 V2V Camera & LiDAR CARLA Intermediate Detection $\times$
AVUCB [152] 2022 V2X Camera MS-COCO [153] Early Detection $\times$
@@
VehIdenti [158] 2023 V2X -- CARTERY [158] -- -- $\times$
@@
DRLCP [163] 2020 V2X Camera CIVS [163] Late -- $\times$
@@
CenterCoop [166] 2023 V2I LiDAR DAIR-V2X Intermediate Detection $\times$
@@
FPV-RCNN [168] 2022 V2V LiDAR COMAP [169] Intermediate Detection $\surd$
MRSR [170] 2023 V2X LiDAR V2X-Sim,OPV2V Intermediate Detection $\surd$
GevBEV [171] 2023 V2V LiDAR OPV2V Late Detection,Segmentation $\times$
  • Note: CAS: Cooperative Agent Selection, PIS: Perception Information Selection, AIS: Joint Agent & Information Selection.
In CP systems relying on V2X communication, the effective deployment of diverse sensors and continuous dissemination of information across multiple agents are essential to provide comprehensive traffic awareness. However, disseminating data without appropriate information filtering can result in substantial communication and computational resource usage, along with a high volume of less pertinent data. For example, neighbouring vehicles might redundantly transmit information about the same objects, causing inefficient utilization of communication channels [173]. This excessive network load elevates the risk of important data packets experiencing delays or being lost, potentially compromising CP performance and safety. To actualize the potential of V2X CP, it is essential to prioritize a swift, secure, and reliable flow of information between vehicles and associated entities like road infrastructures, users, and network nodes, especially given the unpredictable and rapidly changing nature of traffic situations [174].
To avoid communication congestion issues, several researchers have focused on reducing shared data volume before transmission, particularly when bandwidth among agents is limited. As summarized in Table IV, existing strategies aim to alleviate the load on the vehicular network while maintaining the performance of the CP system. These strategies fall into three categories:
  • Cooperative agent selection: This category focuses on minimizing communication costs by transmitting perception information over a partially connected agent graph, thus reducing redundancy;
  • Perception information selection: This category addresses the maximization of informativeness, where cooperative agents select a subset of perceptive data crucial to the ego-vehicle, optimizing the data shared;
  • Agent & information selection: This category encompasses a joint consideration of both cooperative agent selection and perception information selection, aiming to balance communication efficiency and data importance.

### v4

@@ II-D Intermediate Collaboration
III Development of Cooperative Perception: From Early Time Exploration till Recent Research
@@
Traditional approaches in CP have typically relied on late collaboration strategies due to the complexity of feature extraction and limited communication capabilities. With the rapid progress of deep learning, intermediate collaboration has gained prominence in multi-agent cooperation, primarily due to its ability to alleviate the burden on communication resources [111]. Nevertheless, it risks introducing extraneous and unnecessary information, particularly in dynamic networking conditions. To address this challenge, optimizing cooperative information and collaborator selection becomes imperative within a CP system. Furthermore, deep representation introduces higher levels of abstraction and implicitness in the spatial interactions between agents, necessitating suitable data alignment and information fusion algorithms. Developing a unified CP framework to address these issues is of significant importance in advancing Cooperative Driving Automation (CDA).
@@ IV-A Framework Overview
The collaboration stage encompasses four primary components, as depicted in Fig. 5: preparation for transmission, V2X communication, cooperative information alignment, and perception information fusion. To ensure efficient communication while preserving the accuracy of the cooperation system, the initial step involves carefully selecting suitable cooperative agents and pertinent perceptual information. Subsequently, each agent shares the CAM and the refined CPM with its peers through V2X communication [114]. Following this, the ego vehicle aligns the acquired perception information with local data through coordinate conversion or pose transformation techniques. Finally, the aggregated data undergoes consolidation via a fusion algorithm, followed by decoding, resulting in a conclusive perception outcome. Note that all CAVs can transmit their perceptual data to a roadside infrastructure via V2X networks. The associated edge computing infrastructure is responsible for aggregating and consolidating this incoming data, which is then distributed to CAVs according to their specific needs. In the subsequent sections, we provide detailed insights and a review of existing works on the collaboration stage’s primary components.
TABLE III: Summary of the Methods of Preparation for Transmission.
@@
CAS Who2com [115] 2020 -- Camera AirSim-CP Intermediate Segmentation $\surd$
When2com [116] 2020 -- Camera AirSim-MAP Intermediate Segmentation $\surd$
CPOB-LearnCom [117] 2022 V2I LiDAR CARLA-3D Intermediate Detection $\times$
AVUCB [118] 2022 V2X Camera MS-COCO Early Detection $\times$
@@
InterCoop [122] 2024 V2V LiDAR CARLA Intermediate Decision-making $\times$
@@
VehIdenti [124] 2023 V2X -- CARTERY -- -- $\times$
@@
BM2CP [129] 2023 V2V LiDAR & Camera OPV2V, DAIR-V2X Intermediate Detection $\surd$
DRLCP [130] 2020 V2X Camera CIVS Late -- $\times$
FPV-RCNN [131] 2022 V2V LiDAR COMAP Intermediate Detection $\surd$
UMC [132] 2023 V2X LiDAR V2X-Sim,OPV2V Intermediate Detection $\surd$
GevBEV [133] 2023 V2V LiDAR OPV2V Late Detection, Segmentation $\times$
IFTR [134] 2024 V2X Camera DAIR-V2X, V2XSet, OPV2V Intermediate Detection $\surd$
CodeFilling [135] 2024 V2X Camera & LiDAR DAIR-V2X, OPV2VH+ Intermediate Detection $\surd$
V2X-PC [136] 2024 V2X LiDAR DAIR-V2X, V2XSet Early Detection $\times$
PAE [137] 2024 V2X LiDAR CARTI Intermediate Detection $\times$
ActFormer [138] 2024 V2X Camera V2X-Sim Intermediate Detection $\surd$
PIC V2VNet [140] 2020 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\times$
CODS [141] 2023 V2V Camera & LiDAR CARLA Intermediate Detection $\times$
CPSC [142] 2024 V2X LiDAR DAIR-V2X, OPV2V Intermediate Detection $\times$
@@
DiscoNet [144] 2021 V2V LiDAR V2X-Sim Intermediate Detection $\surd$
V2X-ViT [145] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
@@
CenterCoop [147] 2024 V2I LiDAR DAIR-V2X Intermediate Detection $\times$
@@
EMIFF [149] 2024 V2X Camera DAIR-V2X Intermediate Detection $\times$
What2comm [150] 2023 V2X LiDAR DAIR-V2X,V2XSet, OPV2V Intermediate Detection $\times$
SmartCooper [151] 2024 V2V Camera OpenCOOD Intermediate Detection $\times$
  • Note: CAS: Cooperative Agent Selection, PIS: Perception Information Selection, AIS: Agent and Information Selection, PIC: Perception Information Compression.
IV-B Preparation for Transmission
In CP systems, continuous information exchange among multiple agents is crucial for achieving comprehensive traffic awareness. However, real-world application scenarios often face limitations in communication bandwidth. Excessive network load increases the risk of important data packets being delayed or lost, which could compromise the performance and safety of CP systems. Therefore, it is essential to balance the trade-off between perception performance and communication efficiency. To address this challenge, researchers have proposed various strategies aimed at reducing the volume of shared data prior to transmission and alleviating the burden on vehicular networks. As summarized in Table III, these strategies can be grouped into four main approaches:
  • Cooperative Agent Selection: This approach focuses on minimizing communication costs by transmitting perception data through a partially connected agent graph, thereby reducing redundancy within the network.
  • Perception Information Selection: This category aims to select a subset of perceptive data critical to the ego vehicle’s situational awareness. It focuses on maximizing informativeness, which refers to the quality of information, specifically the extent to which the data is relevant, valuable, and reliable in enhancing perception accuracy.
  • Agent and Information Selection: This category encompasses a joint consideration of both cooperative agent selection and perception information selection, aiming to balance communication efficiency and data importance.
  • Perception Information Compression: This strategy involves compressing or encoding essential information before sharing, ensuring that the transmission rate remains within the required data rate limits.